# Latent Denoising Makes Good Visual Tokenizers

### v1

@@ Abstract
Despite their fundamental role, it remains unclear what properties could make visual tokenizers more effective for generative modeling. We observe that modern generative models share a conceptually similar training objective—reconstructing clean signals from corrupted inputs such as Gaussian noise or masking—a process we term denoising. Motivated by this insight, we propose aligning tokenizer embeddings directly with the downstream denoising objective, encouraging latent embeddings to be more easily reconstructed even when heavily corrupted. To achieve this, we introduce the Latent Denoising Tokenizer ($l$-DeTok), a simple yet effective tokenizer trained to reconstruct clean images from latent embeddings corrupted by interpolative noise and random masking. Extensive experiments on ImageNet $256\times 256$ demonstrate that our tokenizer consistently outperforms standard tokenizers across six representative generative models. Our findings highlight denoising as a fundamental design principle for tokenizer development, and we hope it could motivate new perspectives for future tokenizer design. Code is available at: https://github.com/Jiawei-Yang/DeTok
@@ 1 Introduction
Modern visual generative models commonly operate on compact latent embeddings, produced by tokenizers, to circumvent the prohibitive complexity of pixel-level modeling [51, 46, 5, 36]. Current tokenizers are typically trained as standard variational autoencoders [33], primarily optimizing for pixel-level reconstruction. Despite their critical influence on downstream generative quality, it remains unclear what properties enable more effective tokenizers for generation. As a result, tokenizer development has lagged behind recent rapid advances in generative model architectures.
In this work, we ask: What properties can make visual tokenizers more effective for generative modeling? We observe that modern generative models, despite methodological differences, share a conceptually similar training objective—reconstructing original signals from corrupted ones. For instance, diffusion models remove diffusion-induced noise to recover clean signals [31, 46], while autoregressive models reconstruct complete sequences from partially observed contexts [36, 5], analogous to removing “masking noise” [30, 11, 15]. We collectively refer to these reconstruction-from-deconstruction processes as denoising.
@@
Motivated by this insight, we propose to train tokenizers as latent denoising autoencoders, termed as $l$-DeTok. Specifically, we deconstruct latent embeddings via interpolative noise, obtaining corrupted latents by interpolating original embeddings with Gaussian noise. The tokenizer decoder is then trained to reconstruct clean images from these heavily noised latent embeddings. Additionally, we explore random masking, similar to masked autoencoders (MAE) [30], as an alternative form of deconstruction and find it similarly effective.
Conceptually, these deconstruction-reconstruction strategies encourage latent embeddings to be robust, stable, and easily reconstructable under strong corruption, aligning with the downstream denoising tasks central to generative models. Indeed, our experiments show that stronger noise (i.e., strong latent noise and high masking ratio) used in our $l$-DeTok training usually leads to better downstream generative performance.
We demonstrate the effectiveness and generalizability of $l$-DeTok across six representative generative models, including non-autoregressive (DiT [46], SiT [42], LightningDiT [64]) and autoregressive models (MAR [36], RasterAR, RandomAR [36, 44, 68]) on the ImageNet generation benchmark. We find that compared to recent semantics-distilled tokenizers [64, 6], which achieve strong performance on diffusion models but generalize poorly to autoregressive ones, our tokenizer demonstrates significantly broader generalizability. For example, by adopting our tokenizer—without modifying model architectures—we push the limits for MAR models [36], improving FID from 2.31 to 1.55 for MAR-B, matching the performance of the original huge-sized MAR (1.55). For MAR-L, FID improves from 1.78 to 1.35. Importantly, these gains come without semantics distillation, thus avoiding dependencies on visual encoders pretrained at a far larger scale [43, 49].
@@ 2 Related Work
Representation learning in visual recognition.
Representation learning has been a decades-long pursuit in visual recognition, aiming to discover transferable embeddings that generalize across diverse downstream tasks [2]. Starting from supervised representations [17, 24, 65], self-supervised methods have significantly pushed the boundaries of transferability without explicit human annotations [29, 9, 27, 30, 43, 11, 4]. At the core of these approaches lies a foundational principle: pre-training should encourage representations to encode the information most relevant to downstream tasks. This principle has inspired the design of diverse and effective pretext tasks, such as instance discrimination [29, 9, 3], self-distillation [27, 10, 4, 43], and masked-image reconstruction [30, 43, 72], each aligning representations with downstream utility. These insights motivate us to explore tokenizer embeddings that align with downstream generative tasks.
Visual tokenizers for generative modeling.
Modern generative models typically rely on tokenizers to encode images into compact latent embeddings, significantly reducing computational complexity compared to pixel-level modeling [19, 51, 46, 34, 55, 69, 37, 7, 1, 12, 48]. While conventional tokenizers optimize pixel reconstruction with KL-regularization [35], recent approaches [6, 64, 7, 37] have advocated semantics distillation from powerful pretrained vision models [43, 49]. However, such semantics-distilled tokenizers rely inherently on a two-stage pipeline: first training a vision encoder at a significantly larger scale in terms of compute and data, then distilling its features into latent embeddings. Moreover, our experiments reveal a surprising limitation (Table 1): semantics-distilled continuous tokens enhance non-autoregressive diffusion model generation but notably degrade autoregressive methods, questioning the common assumption that tokenizer improvements in one generative paradigm transfer to others. In contrast, our $l$-DeTok generalizes better, consistently benefiting diverse models—yet without relying on semantics distillation.
Generative modeling frameworks.
Generative modeling encompasses various downstream generation frameworks and can be broadly categorized into autoregressive (AR) [8, 19, 5, 36, 59, 21, 44, 68] and diffusion-based non-autoregressive (non-AR) [31, 51, 46, 42, 64] methods. AR models factorize latent representations sequentially, predicting tokens step-by-step conditioned on partially generated contexts, while non-AR models jointly predict all latent tokens through iterative refinement, typically leveraging diffusion [31, 56] or flow-based processes [39, 20]. Despite their methodological diversity, these generative paradigms depend on the quality of latent embeddings produced by upstream tokenizers [28]. This inherent dependence highlights the critical importance of studying latent embeddings: improvements at the tokenizer level have the potential to yield better generative performance across the full spectrum of downstream generative frameworks.
@@ 3.1 Preliminaries of Generative Modeling
Non-autoregressive generative models.
Non-autoregressive models, exemplified by diffusion [31, 46] and flow-matching methods [39, 42], learn to iteratively refine latent representations deconstructed by controlled noise. Given a latent representation of an image $\mathbf{X}_{0}$, the forward noising process progressively corrupts these latents into $\mathbf{X}_{t}$:
$\mathbf{X}_{t}=a(t)\,\mathbf{X}_{0}+b(t)\,\boldsymbol{\epsilon}_{t},\quad\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (1)
@@
$\mathcal{L}(\boldsymbol{\theta})={{\mathbb{E}_{\mathbf{X},\boldsymbol{\epsilon},t}}{\left[{\left\|{{\epsilon_{\boldsymbol{\theta}}}{\left(\mathbf{X}_{t},t\right)}}-\boldsymbol{\epsilon}_{t}\right\|}^{2}\right]}},$ (2)
where ${\epsilon_{\boldsymbol{\theta}}}$ is a learnable noise estimator parameterized by $\boldsymbol{\theta}$. Essentially, non-AR diffusion models learn to reconstruct original latents from intermediate latents deconstructed by noise.
Autoregressive generative models.
Autoregressive approaches factorize image generation into a sequential prediction problem. Given an ordered sequence of latent tokens $\{\mathbf{x}^{1},\dots,\mathbf{x}^{N}\}$, AR methods factorize the joint distribution as:
$p_{\boldsymbol{\theta}}(\mathbf{x})=\prod_{i=1}^{N}p_{\boldsymbol{\theta}}(\mathbf{x}^{i}|\mathbf{x}^{1},\dots,\mathbf{x}^{i-1}),$ (3)
where $\mathbf{x}^{i}$ denotes the latent tokens generated at step $i$. Recent generalized AR variants extend this framework to arbitrary generation orders [36, 68, 44] or set-wise prediction strategies [59, 50]. Nonetheless, the fundamental training objective remains consistent: reconstructing full sequences from partially observed—or equivalently, partially masked—contexts. In other words, AR models learn to reconstruct original latents from intermediate latents deconstructed by masking.
@@ 3.2 Latent Denoising Tokenizers
Overview.
Our tokenizer follows an encoder-decoder architecture based on Vision Transformers (ViT) [18, 61]. Input images are divided into non-overlapping patches, linearly projected into embedding vectors, and added with positional embeddings. During training, we deconstruct these embeddings using two complementary strategies: (i) injecting noise in latent embeddings and (ii) randomly masking image patches. The decoder reconstructs original images from these deconstructed embeddings. This strategy encourages easy-to-reconstruct latent embeddings under heavy corruption, aiming to simplify downstream denoising tasks in generative models.
Noising as deconstruction.
Our core idea is to deconstruct latent embeddings by noise interpolation. Specifically, given latent embeddings $\mathbf{x}$ from the encoder, we interpolate them with Gaussian noise as follows:
$\mathbf{x}^{\prime}=(1-\tau)\mathbf{x}+\tau\boldsymbol{\varepsilon}(\gamma),\quad\text{where}\quad\boldsymbol{\varepsilon}(\gamma)\sim\gamma\cdot\mathcal{N}(\mathbf{0},\mathbf{I}),\quad\tau\sim\mathcal{U}(0,1).$ (4)
Here, the scalar $\gamma$ controls noise standard deviation, and the factor $\tau$ specifies noise level. Critically, this interpolative strategy differs from the conventional additive noise, i.e., $\mathbf{x}^{\prime}=\mathbf{x}+\tau\varepsilon$, employed by standard VAEs [47] or DAEs [62], as it ensures latents can be effectively and heavily corrupted when the noise level $\tau$ is high. Moreover, random sampling of $\tau$ encourages latents to remain robust across diverse corruption levels. Unlike traditional DAEs [62], which apply additive noise directly in pixel space, our latent-space interpolation directly aligns with downstream generative models that operate in latent embedding spaces. At inference time, latent noising is disabled ($\tau=0$).
Masking as deconstruction.
We further generalize our denoising perspective by interpreting masking as another form of latent deconstruction [15, 11, 30]. Inspired by masked autoencoders (MAE) [30], we randomly mask a subset of image patches. Different from MAE, we use a random masking ratio. Concretely, given an input image partitioned into patches, we mask a random subset, where the masking ratio $m$ is sampled from a slightly biased uniform distribution:
@@
where $\mathcal{U}(-0.1,M)$ denotes a uniform distribution on $[-0.1,M]$. The slight bias towards zero reduces the distribution gap between training and inference (no masking). The encoder processes only the visible patches, and masked positions are represented by shared learnable [MASK] tokens at the decoder input. At inference time, all patches are visible ($m=0$).
Training objectives.
Our decoder reconstructs the original images from corrupted latent embeddings. The training objective follows established practice [51, 19, 69], combining pixel-wise mean-squared-error (MSE), latent-space KL-regularization [47], perceptual losses (VGG- [54] and ConvNeXt-based [40] as in [69]), and an adversarial GAN objective [25]:
@@ 4 Implementation
We describe the implementation details, including datasets, evaluation metrics, tokenizer and generative model training procedures.
Dataset and metrics.
We perform all experiments on ImageNet [52] at $256\times 256$ resolution. Images are center-cropped, randomly horizontally flipped during training, and normalized to $[-1,1]$. We evaluate generative models using Fréchet Inception Distance (FID), Inception Score (IS), precision and recall, following [16]. For tokenizer reconstruction, we report reconstruction-FID (rFID).
Tokenizer baselines.
@@
We benchmark our tokenizer against a diverse set of publicly available tokenizers: (i) MAR-VAE from MAR [36], trained on ImageNet using the implementation from [19]; (ii) VA-VAE [64], aligning latent embeddings with DINOv2 features [43]; (iii) MAETok [6], distilling HOG [13], DINOv2 [43], and CLIP [49] features through auxiliary decoders; (iv) SD-VAE from Stable-Diffusion [51], trained on significantly larger datasets. Additionally, for controlled comparisons, we also train our own baseline tokenizer without the proposed denoising approaches.
Our tokenizer implementation.
@@
We implement our tokenizer using ViTs for both encoder and decoder [61, 18]. We adopt recent architectural advances from LLaMA [60], including RoPE [57] (together with learned positional embeddings following [22]), RMSNorm [71], and SwiGLU-FFN [53]. The encoder operates at a patch size of $16$, yielding $256$ latent tokens for each $256\times 256$ image, while the decoder uses patch size $1$ as there is no resolution change. The latent dimension is set to $16$. We omit the [CLS] token for simplicity.
Tokenizer training. Our tokenizer is trained using the weighted loss defined in Eq. 6, with default weights $\lambda_{\text{KL}}=10^{-6}$, $\lambda_{\text{percep}}=1.0$, and $\lambda_{\text{GAN}}=0.1$. In ablation studies, we use ViT-S for the encoder and ViT-B for the decoder, disable the GAN loss, and train for 50 epochs. We observe that including a GAN loss [66, 25, 19] sharpens reconstructions but roughly doubles training time, without altering result trends. For final experiments, we use ViT-B for both encoder and decoder, train for 200 epochs, and activate the GAN loss starting from epoch 100. All tokenizers use AdamW [41] with a global batch size of 1024 and peak learning rate of $4.0\times 10^{-4}$ (corresponding to a base learning rate of $1.0\times 10^{-4}$ scaled linearly by the global batch size divided by 256 [26]), linear warm-up, and cosine learning schedule. Further implementation details are provided in the Appendix.
Generative models.
@@
To evaluate the broad effectiveness of a tokenizer, we experiment with six representative generative models, including three non-autoregressive models: DiT [46], SiT [42], and LightningDiT [64]; and three autoregressive models: MAR [36], causal RandomAR, and RasterAR based on RAR [68] and diffloss [36]. We follow the officially released implementations to reimplement all methods within a unified codebase, standardizing training and evaluation across methods. Previous works on visual tokenizers often exclusively evaluate on non-AR models (e.g., DiT, SiT), yet we find improvements from non-AR models do not necessarily translate to AR models (more on this later). Our experiments aim to provide useful data points toward a more universal tokenizer.
Generative model training.
@@
We use a standardized training recipe for all generative models. Specifically, we follow the hyperparameters from [64], training generative models with a global batch size of 1024, using AdamW [41] with a constant learning rate of $2\times 10^{-4}$, without warm-up, gradient clipping, or weight decay. Autoregressive models adopt the three-layer 1024-channel diffloss MLP from [36]. For ablation studies, we train generative models for 100 epochs. For larger-scale experiments on MAR models, we train them for 800 epochs. All models utilize exponential moving average (EMA) with a decay rate of 0.9999. We always standardize tokenizer outputs by subtracting the mean and dividing by the standard deviation, both computed from the ImageNet training set. For publicly available tokenizers, we use their official standardization steps. Unless noted otherwise, we report FID@50k scores with classifier-free guidance (CFG), with optimal CFG scales searched from FID@10k results. We abbreviate FID@50k as FID.

### v2

# Latent Denoising Makes Good Tokenizers

@@ Abstract
Despite their fundamental role, it remains unclear what properties could make tokenizers more effective for generative modeling. We observe that modern generative models share a conceptually similar training objective—reconstructing clean signals from corrupted inputs, such as signals degraded by Gaussian noise or masking—a process we term denoising. Motivated by this insight, we propose aligning tokenizer embeddings directly with the downstream denoising objective, encouraging latent embeddings that remain reconstructable even under significant corruption. To achieve this, we introduce the Latent Denoising Tokenizer ($l$-DeTok), a simple yet highly effective tokenizer trained to reconstruct clean images from latent embeddings corrupted via interpolative noise or random masking. Extensive experiments on class-conditioned (ImageNet $256\times 256$ and $512\times 512$) and text-conditioned (MSCOCO) image generation benchmarks demonstrate that our $l$-DeTok consistently improves generation quality across six representative generative models compared to prior tokenizers. Our findings highlight denoising as a fundamental design principle for tokenizer development, and we hope it could motivate new perspectives for future tokenizer design. Code is available at: https://github.com/Jiawei-Yang/DeTok.
@@ 1 Introduction
Modern visual generative models commonly operate on compact latent embeddings, produced by tokenizers, to circumvent the prohibitive complexity of pixel-level modeling (Rombach et al., 2022; Peebles and Xie, 2023; Chang et al., 2022; Li et al., 2024a). Current tokenizers are typically trained as standard variational autoencoders (Kingma and Welling, 2014), primarily optimizing for pixel-level reconstruction. Despite their critical influence on downstream generative quality, it remains unclear what properties enable more effective tokenizers for generation. As a result, tokenizer development has lagged behind recent rapid advances in generative model architectures.
In this work, we ask: What properties can make visual tokenizers more effective for generative modeling? We observe that modern generative models, despite methodological differences, share a conceptually similar training objective—reconstructing original signals from corrupted ones. For instance, diffusion models remove diffusion-induced noise to recover clean signals (Ho et al., 2020; Peebles and Xie, 2023), while autoregressive models reconstruct complete sequences from partially observed contexts (Li et al., 2024a; Chang et al., 2022), analogous to removing “masking noise” (He et al., 2022; Chen et al., 2025c; Devlin et al., 2019). We collectively refer to these reconstruction-from-deconstruction [^1] We use the terms “deconstruction” and “corruption” interchangeably throughout the paper. processes as denoising.
@@
Motivated by this insight, we propose to train tokenizers as latent denoising autoencoders, termed as $l$-DeTok. Specifically, we corrupt latent embeddings via interpolative noise, generated by interpolating original embeddings with Gaussian noise. The tokenizer decoder is then trained to reconstruct clean images from these heavily noised latent embeddings. Additionally, we explore random masking, akin to masked autoencoders (MAE) (He et al., 2022), as an alternative and optional form of deconstruction and find it similarly effective.
Conceptually, these deconstruction-reconstruction strategies encourage latent embeddings to be robust, stable, and easily reconstructable under strong corruption, aligning with the downstream denoising tasks central to generative models. Indeed, our experiments show that stronger noise used in our $l$-DeTok training typically yields better downstream generative performance.
We demonstrate the effectiveness and generalizability of $l$-DeTok across six representative generative models, including non-autoregressive (DiT (Peebles and Xie, 2023), SiT (Ma et al., 2024), LightningDiT (Yao and Wang, 2025)) and autoregressive models (MAR (Li et al., 2024a), RasterAR, RandomAR (Li et al., 2024a; Pang et al., 2025; Yu et al., 2024)) on the ImageNet generation benchmark. By adopting our tokenizer—without modifying model architectures—we push the limits for MAR models (Li et al., 2024a), improving FID from 2.31 to 1.55 for MAR-B, matching the performance of the original huge-sized MAR (1.55). For MAR-L, FID improves from 1.78 to 1.35. Importantly, these gains come without semantics distillation, thus avoiding reliance on visual encoders pretrained at a far larger scale (Oquab et al., 2024; Radford et al., 2021).
@@ 2 Related Work
Representation learning in visual recognition. Representation learning has been a decades-long pursuit in visual recognition, aiming to obtain transferable embeddings that generalize across diverse downstream tasks (Bengio et al., 2013). At the core of these approaches lies a foundational principle: pre-training should encourage representations to encode the information most relevant to downstream tasks. This principle has inspired the design of diverse and effective pretext tasks, such as instance discrimination (He et al., 2020; Chen et al., 2020b), self-distillation (Grill et al., 2020; Caron et al., 2021; Oquab et al., 2024), and masked-image reconstruction (He et al., 2022; Oquab et al., 2024; Zhou et al., 2022), each aligning representations with downstream utility. These insights motivate us to explore tokenizer embeddings that align with downstream generative tasks.
Visual tokenizers for generative modeling. Modern generative models rely on tokenizers to encode images into compact latent embeddings, significantly reducing computational complexity compared to pixel-level modeling. While conventional tokenizers optimize pixel reconstruction with KL-regularization, recent approaches (Chen et al., 2025a; Yao and Wang, 2025; Chen et al., 2025b; Li et al., 2025) have advocated semantics distillation from powerful pretrained vision models (Oquab et al., 2024; Radford et al., 2021). Yet, in many domains (e.g., video, audio, 3D/4D), such encoders may not exist. In contrast, our $l$-DeTok learns good tokenizers without these dependencies, and it could be complementary to approaches such as $\epsilon$-VAE, which replaces the deterministic VAE decoder with a denoising diffusion process guided by encoder latents (Zhao et al., 2025), and residual vector-quantized tokens, which support high-fidelity yet efficient discrete diffusion generation via collective token prediction (Kim et al., 2025).
Generative modeling frameworks. Generative modeling frameworks broadly include autoregressive (AR) (Chen et al., 2020a; Esser et al., 2021; Chang et al., 2022; Li et al., 2024a; Tian et al., 2025; Pang et al., 2025; Yu et al., 2024; Tschannen et al., 2024) and diffusion-based non-autoregressive (non-AR) methods (Ho et al., 2020; Rombach et al., 2022; Peebles and Xie, 2023; Ma et al., 2024; Yao and Wang, 2025). AR models predict latent tokens sequentially conditioned on partial contexts, while non-AR models jointly generate tokens via iterative refinement, typically using diffusion (Ho et al., 2020; Sohl-Dickstein et al., 2015) or flow-based processes (Lipman et al., 2023; Esser et al., 2024). Despite methodological differences, both paradigms critically rely on high-quality latent embeddings from upstream tokenizers (Hansen-Estruch et al., 2025). Thus, improving tokenizer embeddings is fundamental for enhancing generative performance across diverse frameworks.
@@ 3.1 Preliminaries of Generative Modeling
Non-autoregressive generative models. Non-autoregressive models, exemplified by diffusion (Ho et al., 2020; Peebles and Xie, 2023) and flow-matching methods (Lipman et al., 2023), learn to iteratively refine latent representations deconstructed by controlled noise. Given a latent representation of input $\mathbf{X}_{0}$, the forward noising process progressively corrupts these latents into $\mathbf{X}_{t}$:
$\mathbf{X}_{t}=a(t)\,\mathbf{X}_{0}+b(t)\,\bm{\epsilon}_{t},\quad\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
@@
$\mathcal{L}(\bm{\theta})={{\mathbb{E}_{\mathbf{X},\bm{\epsilon},t}}{\left[{\left\|{{\epsilon_{\bm{\theta}}}{\left(\mathbf{X}_{t},t\right)}}-\bm{\epsilon}_{t}\right\|}^{2}\right]}},$ (2)
where $\epsilon_{\bm{\theta}}$} is a learnable noise estimator parameterized by $\bm{\theta}$. Essentially, non-AR diffusion models learn to reconstruct original latents from intermediate latents deconstructed by noise.
Autoregressive generative models. Autoregressive approaches factorize image generation into a sequential prediction problem. Given an ordered sequence of latent tokens $\{\mathbf{x}^{1},\dots,\mathbf{x}^{N}\}$, AR methods factorize the joint distribution as:
$p_{\bm{\theta}}(\mathbf{x})=\prod_{i=1}^{N}p_{\bm{\theta}}(\mathbf{x}^{i}|\mathbf{x}^{1},\dots,\mathbf{x}^{i-1}),$ (3)
where $\mathbf{x}^{i}$ denotes the latent tokens generated at step $i$. Recent generalized AR variants extend this framework to arbitrary generation orders (Li et al., 2024a; Yu et al., 2024; Pang et al., 2025) or set-wise prediction strategies (Tian et al., 2025; Ren et al., 2025). Nonetheless, the fundamental training objective remains consistent: reconstructing full sequences from partially observed—or equivalently, partially masked—contexts. In other words, AR models learn to reconstruct original latents from intermediate latents deconstructed by masking.
@@ 3.2 Latent Denoising Tokenizers
Overview. Our tokenizer follows an encoder-decoder architecture based on Vision Transformers (ViT) (Dosovitskiy et al., 2021). Input images are divided into non-overlapping patches, linearly projected into embedding vectors, and added with positional embeddings. During training, we deconstruct these embeddings using two complementary strategies: (i) injecting noise in latent embeddings and (ii) randomly masking image patches. The decoder reconstructs original images from these deconstructed embeddings. This strategy encourages easy-to-reconstruct latent embeddings under heavy corruption, aiming to simplify downstream denoising tasks in generative models.
Noising as deconstruction. Our core idea is to deconstruct latent embeddings by noise. Specifically, given latent embeddings $\mathbf{x}$ from the encoder, we interpolate them with Gaussian noise as follows:
$\mathbf{x}^{\prime}=(1-\tau)\mathbf{x}+\tau\bm{\varepsilon}(\gamma),\quad\text{where}\quad\bm{\varepsilon}(\gamma)\sim\gamma\cdot\mathcal{N}(\mathbf{0},\mathbf{I}),\quad\tau\sim\mathcal{U}(0,1).$ (4)
Here, the scalar $\gamma$ controls noise standard deviation, and the factor $\tau$ specifies noise level. Critically, this interpolative strategy differs from the conventional additive noise, i.e., $\mathbf{x}^{\prime}=\mathbf{x}+\tau\bm{\varepsilon}$, employed by standard VAEs (Pu et al., 2016), as it ensures latents can be effectively and heavily corrupted when the noise level $\tau$ is high. Moreover, random sampling of $\tau$ encourages latents to remain robust across diverse corruption levels. At inference time, latent noising is disabled ($\tau=0$).
Masking as deconstruction. We further generalize our denoising perspective by interpreting masking as another form of latent deconstruction (Chen et al., 2025c). Inspired by masked autoencoders (MAE) (He et al., 2022), we randomly mask a subset of image patches. Different from MAE, we use a random masking ratio. Concretely, given an input image partitioned into patches, we mask a random subset, where the masking ratio $m$ is sampled from a slightly biased uniform distribution:
@@
where $\mathcal{U}(-0.1,M)$ denotes a uniform distribution on [0.1,M][-0.1,M]. The slight bias towards zero reduces the distribution gap between training and inference (no masking). The encoder processes only the visible patches, and masked positions are represented by shared learnable [MASK] tokens at the decoder input. At inference time, all patches are visible ($m=0$).
Training objectives. Our tokenizer is trained to reconstruct the original images from corrupted latent embeddings. The training objective follows established practice (Rombach et al., 2022; Esser et al., 2021; Yu et al., 2025a), combining pixel-wise mean-squared-error (MSE), latent-space KL-regularization (Pu et al., 2016), perceptual losses (VGG- and ConvNeXt-based as in Yu et al. (2025a)), and an adversarial GAN objective (Goodfellow et al., 2014):
@@ 4 Implementation
We briefly summarize implementation details here, including datasets, evaluation metrics, and model training procedures. Due to limited space, we defer comprehensive details to Sections A.1 and A.2.
Dataset and metrics. We primarily experiment with class-conditioned image generation on the ImageNet dataset (Russakovsky et al., 2015) at $256\times 256$ and $512\times 512$ resolutions, and further evaluate text-to-image generation on the MS-COCO dataset (Lin et al., 2014). For evaluation, we report Fréchet Inception Distance (FID), Inception Score (IS), precision, and recall.
@@
Tokenizer baselines. We benchmark our tokenizer against a diverse set of tokenizers: (i) MAR-VAE from MAR (Li et al., 2024a), trained on ImageNet using the implementation from Esser et al. (2021); (ii) VA-VAE (Yao and Wang, 2025), aligning latent embeddings with DINOv2 features; (iii) MAETok (Chen et al., 2025a), distilling HOG, DINOv2 (Oquab et al., 2024), and CLIP (Radford et al., 2021) features through auxiliary decoders; (iv) SD-VAE and SD3-VAE from Stable-Diffusion (Rombach et al., 2022; Esser et al., 2024), trained on significantly larger datasets. Additionally, for controlled comparisons, we also train our ViT-based tokenizer without denoising objective as a baseline.
@@
Our tokenizer implementation. We implement our tokenizer using ViTs for both encoder and decoder (Vaswani et al., 2017; Dosovitskiy et al., 2021). We adopt recent architectural advances from LLaMA (Touvron et al., 2023), including RoPE (Su et al., 2024) (together with learned positional embeddings following Fang et al. (2023)), RMSNorm (Zhang and Sennrich, 2019), and SwiGLU-FFN (Shazeer, 2020). We use a patch size of $16$, and the latent dimension is set to $16$.
Tokenizer training. We set the default weights in Eq. 6 to $\lambda_{\text{KL}}=10^{-6}$, $\lambda_{\text{percep}}=1.0$, and $\lambda_{\text{GAN}}=0.1$. In ablation studies, we use ViT-S for the encoder and ViT-B for the decoder, disable the GAN loss, and train for 50 epochs. For final experiments, we use ViT-B for both encoder and decoder, train for 200 epochs, and activate the GAN loss starting from epoch 100.
@@
Generative models. To evaluate the broad effectiveness of a tokenizer, we experiment with six representative generative models, including three non-autoregressive models: DiT (Peebles and Xie, 2023), SiT (Ma et al., 2024), and LightningDiT (Yao and Wang, 2025); and three autoregressive models: MAR (Li et al., 2024a), causal RandomAR, and RasterAR based on RAR (Yu et al., 2024) and diffloss (Li et al., 2024a). We follow the officially released implementations to reimplement all methods within a unified codebase, standardizing training and evaluation across methods. Previous works on visual tokenizers often exclusively evaluate on non-AR models (e.g., DiT, SiT), yet we find improvements from non-AR models do not necessarily generalize to AR models (more on this later). Our experiments aim to provide useful data points toward a more universal tokenizer.
@@
Generative model training. We use a standardized training recipe for all generative models. Specifically, we follow the hyperparameters from Yao and Wang (2025), training generative models with a global batch size of 1024, using AdamW (Loshchilov and Hutter, 2019) with a constant learning rate of $2\times 10^{-4}$, without warm-up, gradient clipping, or weight decay. Autoregressive models adopt the three-layer 1024-channel diffloss MLP from Li et al. (2024a). For ablation studies, we train generative models for 100 epochs. For larger-scale experiments on MAR models, we train for 800 epochs.
Figure 2: Ablation on latent noise design. (a) Additive vs. interpolative noise. Interpolative noise clearly outperforms additive noise for both MAR (Li et al., 2024a) and SiT (Ma et al., 2024). Both interpolative and additive latent noise lead to improved performance for MAR. (b) Latent noise standard deviation ($\gamma$). Our $l$-DeTok remains robust across various noise standard deviations. Generally, increasing $\gamma$ improves generation quality, with best results achieved around $\gamma=3.0$.